# Extreme Jet Beaming Observed in Neutrino-Associated Blazars

### v1

@@ 1 Introduction
Blazars were the first high-energy (TeVs-PeVs) astrophysical neutrino sources to be discovered (IceCube Collaboration et al., 2018). Later, neutrinos have also been detected from other sources, including the Milky Way Galaxy (Kovalev et al., 2022; Icecube Collaboration et al., 2023; Albert et al., 2023; Allakhverdyan et al., 2024a), nearby Seyfert-type Active Galactic Nuclei (AGNs, IceCube Collaboration et al., 2022; Neronov et al., 2024), and Tidal Disruption Events in AGNs (TDEs, van Velzen et al., 2024). Still, blazars remain a very prominent class of bright neutrino emitters. There are more and more population-based studies resulting in neutrinos associated with AGNs selected by their compact radio flux (basically, blazars; e.g., P23; Kouch et al. 2024; Albert et al. 2024; Suray & Troitsky 2024), X-ray flux (Kun et al. 2024; Khatee Zathul et al. 2024; Neronov et al. 2024; blazars in Plavin et al. 2024), or using more specialized blazar selection methods (e.g., Kadler et al., 2016; Giommi et al., 2020; Buson et al., 2022).
Recent studies (e.g. P23; Kouch et al. 2024) have specifically demonstrated a strong observational connection of radio-bright blazars being also bright neutrino sources. This connection is manifested both as the neutrino – pc-scale radio flux correlation (see Section 3 for updates), and also as a temporal correlation with neutrinos coming predominantly during radio flares. Those results already provide an indication that neutrinos are emitted predominantly along the jet axis in blazars; more direct studies should help quantifying this effect. Furthermore, note that the strong neutrino-blazars association is not found in every analysis, depending on the handling of directional uncertainties (see discussion in Abbasi et al., 2023; Kouch et al., 2024). Utilizing additional observables is thus crucial to refine and potentially confirm the connection.
@@ 2 Data
Figure 1: The sky distribution of radio blazars and IceCube events used in this work. See Section 2 for the definitions of each sample. Two different IceCube event selections are shown in different manners (filled and unfilled) to enable a clearer visual comparison between them. There are 83 blazars from the RFC catalog coincident with at least one IceCube neutrino; note that they include both real physical associations and chance coincidences, see Figure 2. Neutrino-coincident MOJAVE 1.5JyQC (Section 2.1) blazars are labeled here; see Figure 3 for their VLBI images and Table 1 for their properties.
@@ 2.1 Radio observations of parsec-scale emission
For Doppler boosting and jet viewing angle measurements, we rely on the results of the MOJAVE (Monitoring Of Jets in Active galactic nuclei with VLBA Experiments) team. These parameters cannot be measured directly for blazars; the MOJAVE team has suggested (Homan et al., 2006) and applied (Homan et al., 2021) a procedure to infer them from VLBI monitoring data; we use those measurements in this paper. The full MOJAVE dataset is assembled with a range of selection criteria: bright VLBI-selected blazars, compact symmetric objects, Fermi-selected blazars, low-luminosity AGNs (Lister et al., 2011, 2019, 2020, 2013). Similarly to our earlier studies, we focus on a uniform flux-density limited subsample of the MOJAVE program sources in the analysis: the 1.5 Jy Quarter Century Sample (1.5JyQC; defined in Lister et al. 2019). It consists of blazars at $\delta>-30\degree$ with VLBI flux density exceeding 1.5 Jy at 15 GHz at least once during the quarter century of observations, 1994–2019. This sample contains 232 blazars, and 205 of those have Doppler boosting measurements (highlighted in Figure 1). The MOJAVE 1.5JyQC sample was specifically designed to be homogeneous and unbiased.
@@ 2.2 IceCube high-energy neutrinos
The first catalog is the sample of high-energy alert-like neutrino detections we collected before, most recently in P23. Earlier, the same selection criteria were utilized in P20 and Hovatta et al. (2021). This sample consists of 71 events with the angular uncertainty region below 10 sq. deg. and the estimated energy above 200 TeV.
@@
For the analyses performed in this work, we enlarge the original IceCube directional 90% uncertainties by $0.78\degree$ isotropically in all directions. This enlargement was determined in P23 to be optimal for association searches, and we keep the same value to avoid multiple trials in the statistical analysis. A similar enlargement was found in Kouch et al. (2024); see that paper, P23, and Section 3 for a more detailed discussion.
The sky localization of IceCube events from both samples is shown in Figure 1. Blazars coincident with IceCube events are highlighted in the figure, and those from the MOJAVE 1.5JyQC sample are labeled; see Table 1 for their properties.
@@ 3 Neutrino-blazar correlation
2015--09--26 IC150926A J1256--0547 1253--055 $0.536$ $15$ $19$ yes $140$ $72$ $0.12$
2022--03--03 J1751$+$0939 1749$+$096 $0.322$ $2.8$ $4$ yes $100$ $51$ $0.07$
2019--07--30 IC190730A J1504$+$1029 1502$+$106 $1.838$ $1.5$ $1.4$ yes $65$ $35$ $0.44$
2016--01--28 IC160128A J1733--1304 1730--130 $0.902$ $4$ $4.5$ yes $48$ $32$ $1$
2015--08--12 IC150812B J2148$+$0657 2145$+$067 $0.999$ $6.6$ $7.5$ yes $31$ $16$ $0.36$
2015--08--31 IC150831A J0336$+$3218 0333$+$321 $1.259$ $1.6$ $1.7$ yes $30$ $18$ $1.4$
2012--05--15 IC120515A J1310$+$3220 1308$+$326 $0.997$ $1.9$ $2.1$ yes $24$ $28$ $2.3$
2022--02--05 IC220205B J1743--0350 1741--038 $1.054$ $4$ $5.3$ yes $22$
2011--09--30 J1743--0350 1741--038 $1.054$ $4$ $5.3$ yes $22$
2013--10--14 IC131014A J0211$+$1051 0208$+$106 $0.2$ $0.66$ $1.1$ yes $6.3$ $5.1$ $8.8$
2020--11--30 IC201130A J0206--1150 0203--120 $1.663$ $0.28$ $0.3$ no $35$ $20$ $1$
2017--09--22 IC170922A J0505$+$0459 0502$+$049 $0.954$ $0.7$ $0.93$ no $25$ $14$ $1.6$
IC221223A J2311$+$3425 2308$+$341 $1.817$ $0.69$ $0.86$ no $23$
IC201221A J1724$+$4004 1722$+$401 $1.049$ $0.5$ $0.58$ no $9.8$ $61$ $3.2$
2010--10--09 J2200$+$1030 2157$+$102 $0.172$ $0.19$ $0.22$ no $5$
2017--09--22 IC170922A J0509$+$0541 0506$+$056 $0.3365$ $0.42$ $0.46$ no $1.8$ $1.5$ $33$
2011--07--14 IC110714A J0432$+$4138 0429$+$415 $1.022$ $1.4$ $1$ no $0.13$ $54$ $31$
Note. — Columns are as follows: (1) — IceCube neutrino arrival date for events included in the P23 sample; (2) — IceCat event ID for events included in IceCat-1; see Section 2.2 for a definition of these samples; (3), (4) — coordinate-based blazar names, with the J2000 name taken directly from the RFC (Petrov & Kovalev, 2025); (5) — redshift $z$ extracted from the NASA/IPAC Extragalactic Database (NED); (6), (7) — parsec-scale radio flux densities measured by VLBI, as presented in the RFC catalog; their typical uncertainty is about $10\%$, dominated by the amplitude calibration; (8) — flag indicating whether the blazar is included in the uniformly-selected MOJAVE 1.5JyQC sample (Section 2.1); those sources are also labeled in Figure 1, and their VLBI images are shown in Figure 3; (9), (10), (11) — beaming indicators, Doppler $D$ and Lorentz $\Gamma$ factors and the viewing angle $\theta$, as measured within the MOJAVE program (Homan et al., 2021); see Section 4.1 for a discussion of assumptions and uncertainties of these measurements.
@@
After having chosen the test statistic, we calculate the probability that its observed value could have arisen by chance, assuming no directional neutrino-blazar correlation. For that, we randomize the Right Ascensions of all IceCube events $10^{6}$ times, motivated by the instrument sensitivity being dependent on the Declination only (used in Aartsen et al., 2017, and elsewhere). Then, the chance probability is the fraction of random realizations where the average VLBI flux of matching blazars is higher than the observed value. There are no free parameters to be optimized in this procedure, and we can immediately interpret the resulting fraction as the $p$-value.
@@
Note that updated or more detailed event reconstructions are not necessarily better in every aspect: see demonstrations and discussion in the review Troitsky (2024) and in Sommani et al. (2024). For example, a reconstruction update lead to a decrease of the TXS 0506+056 flare significance (IceCube Collaboration et al., 2021; Abbasi et al., 2023). A detailed analysis and interpretation of the differences between various IceCube event reconstructions is out of scope of this work.
@@
As detailed in Section 2.2, we enlarge original IceCube event uncertainties by $0.78\degree$ for all calculations in this paper, following the approach of P23 and Kouch et al. (2024) and avoiding multiple comparisons issues. Here, we present a focused cross-check that such an enlargement remains important for both IceCube event catalogs we use (Section 2.2).
Repeating our statistical testing procedure with original IceCube uncertainties taken as-is, we don’t see any significant evidence for the correlation with VLBI blazars. The chance probability $p>5\%$ for both IceCube event selections. The spatial correlation shown above is heavily driven by blazars that lie slightly outside the original reported 90% directional errors. This is still consistent with the observed correlation caused by a physical connection. Such an effect could result from unaccounted systematic uncertainties or other discrepancies between estimated and actual error regions. A detailed dedicated study can only be performed by the IceCube collaboration, utilizing all the available data about each detection; see Lagunas Gualda et al. (2021) for a discussion and an overview of challenges. Further in the future, it will be crucial to compare the brightest neutrino sources in the sky between different neutrino observatories. The cross-check we performed here does not replace the detailed analysis performed in P23 and Kouch et al. (2024), but remains consistent with those works and results reported by the IceCube (Abbasi et al., 2023, 2024a).
@@ 4 Geometry and kinematics of neutrino-associated blazars
Figure 3: MOJAVE VLBI images for the nine 1.5JyQC blazars coincident with IceCube neutrinos from both samples of Section 2. We show here averages of multi-decade individual VLBI images, which provide a better dynamic range to trace the overall structure of the jet. These stacked images are available from the MOJAVE website (https://www.cv.nrao.edu/MOJAVE/), see details in Pushkarev et al. (2017). For each blazar, we show the redshift $z$, the Doppler factor $D$ and the viewing angle $\theta$ estimates. Blazars are ordered by their Doppler factor, same as in Table 1.
@@ 4.1 Observational evidence for neutrino beaming
The association of high-energy neutrinos with VLBI-bright blazars (Section 3) already suggests that beaming plays a major role in neutrino production and detection: neutrinos are predominantly emitted along the direction of the jet. Indeed, VLBI-selected AGNs are dominated by galaxies with jets pointing towards us, their typical viewing angles are $\theta=2\dots 3\degree$ (Lister et al., 2019; Homan et al., 2021). If neutrino emission were isotropic, we would expect to see thousands ($30004\pi/\pi\theta^{2}~{}\sim~{}$ 3000) of similar AGNs pointed in other directions for each neutrino-VLBI association. This is clearly incompatible with the observed number of high-energy neutrinos.
@@
They are measured from VLBI observations within the MOJAVE program (Section 2.1, Homan et al. 2021). These estimates crucially rely on the VLBI ability to resolve the parsec-scale structure and determine the brightness of the most compact detail – the so-called core. Due to the measurement procedure, the estimates of these quantities are somewhat related to the VLBI flux density we use in the correlation analysis above (Section 3). Still, they are not exactly the same and differ in fundamental ways. The VLBI flux density is a measure of the total radio emission on multi-parsec scale, while Doppler factors are determined from the brightness of even more compact regions, typically a fraction of a parsec in projection. Viewing angle and Lorentz factor estimates additionally rely on parsec-scale kinematics to constrain the geometry. It can be challenging to calculate precise uncertainties of these measurements. Still, as a cross-check, Doppler factors resulting from MOJAVE measurements were found to be consistent with independent estimates obtained via different techniques, such as radio variability (Liodakis et al., 2018; Homan et al., 2021).
@@
We evaluate the statistical significance of the Doppler factor correlation with neutrino coincidences. Doppler factors are chosen because they are the most direct tracers of the electromagnetic emission beaming. They are also measured more directly from observations, see Homan et al. (2021). We rely on the same statistical testing procedure as in Section 3, namely the shuffling of neutrino event positions in the sky. We choose the Kolmogorov-Smirnov test statistic between the Doppler factor distributions for MOJAVE 1.5JyQC blazars and for neutrino coincidences.
@@
The most extreme viewing angle estimates from the MOJAVE observations can be less than $0.1\degree$. There are only three sources with such viewing angles in the full sample (Homan et al., 2021), and one of them is coincident with an IceCube neutrino: 1749+096 with $\theta=0.07\degree$, see Table 1 and Figure 3. We believe that such blazars do indeed have jets pointing almost directly towards us, but recommend treating specific values of these extreme viewing angle estimates with caution. It is possible that some limitations of the generic approach utilized in Homan et al. (2021) manifest in such extreme cases. Additional or alternative ways to infer the jet geometry and the beaming patterns would potentially be useful for more detailed quantitative studies in the future. For example, a very large apparent jet opening angle (Pushkarev et al., 2017) might be used as another independent indicator of a very small viewing angle (e.g., the case of 1424+240, Kovalev et al., in prep.).

### v2

@@ 1 Introduction
Blazars were the first high-energy (TeVs-PeVs) astrophysical neutrino sources to be discovered (IceCube Collaboration et al., 2018). Later, neutrinos have also been detected from other sources, for instance the Milky Way Galaxy (Y. Y. Kovalev et al., 2022; Icecube Collaboration et al., 2023; A. Albert et al., 2023; V. A. Allakhverdyan et al., 2024a) and nearby Seyfert-type Active Galactic Nuclei (AGNs, R. Abbasi et al., 2022; A. Neronov et al., 2024). Still, blazars remain a very prominent class of bright neutrino emitters. There are more and more population-based studies resulting in neutrinos associated with AGNs selected by their compact radio flux (basically, blazars; e.g., A. V. Plavin et al. 2023; P. M. Kouch et al. 2024; A. Albert et al. 2024; A. Suray & S. Troitsky 2024), X-ray flux (E. Kun et al. 2024; A. Neronov et al. 2024; blazars in A. V. Plavin et al. 2024), or using more specialized blazar selection methods (e.g., M. Kadler et al., 2016; P. Giommi et al., 2020; S. Buson et al., 2022). In this work, we adopt the commonly used definition of “blazar”, referring to an AGN with a powerful jet aligned close to our line of sight.
Recent studies (e.g. A. V. Plavin et al. 2023; P. M. Kouch et al. 2024) have specifically demonstrated a strong observational connection of radio-bright blazars being also bright neutrino sources. This connection is manifested both as the neutrino – pc-scale radio flux correlation (see Section 3 for updates), and also as a temporal correlation with neutrinos coming predominantly during radio flares. Those results already provide an indication that neutrinos are emitted predominantly along the jet axis in blazars; more direct studies should help quantifying this effect. Furthermore, note that the strong neutrino-blazars association is not found in every analysis, depending on the handling of directional uncertainties (see discussion in R. Abbasi et al., 2023; P. M. Kouch et al., 2024). Utilizing additional observables is thus crucial to refine and potentially confirm the connection.
@@ 2 Data
Figure 1: The sky distribution of radio blazars and IceCube events used in this work. See Section 2 for the definitions of each sample. Two different IceCube event selections are shown in different manners (filled and unfilled) to enable a clearer visual comparison between them. There are 83 blazars from the RFC catalog coincident with at least one IceCube neutrino; note that they include both real physical associations and chance coincidences, see Figure 2. Neutrino-coincident MOJAVE 1.5JyQC (Section 2.1) blazars are labeled here; see Figure A1 for their VLBI images and Table 1 for their properties.
@@ 2.1 Radio observations of parsec-scale emission
For Doppler boosting and jet viewing angle measurements, we rely on the results of the MOJAVE (Monitoring Of Jets in Active galactic nuclei with VLBA Experiments) team. These parameters cannot be measured directly for blazars; the MOJAVE team has suggested (D. C. Homan et al., 2006) and applied (D. C. Homan et al., 2021) a procedure to infer them from VLBI monitoring data; we use those measurements in this paper. The full MOJAVE dataset is assembled with a range of selection criteria: bright VLBI-selected blazars, compact symmetric objects, Fermi-selected blazars, low-luminosity AGNs (M. L. Lister et al., 2011, 2019, 2020, 2013). Similarly to our earlier studies, we focus on a uniform flux-density limited subsample of the MOJAVE program sources in the analysis: the 1.5 Jy Quarter Century Sample (1.5JyQC; defined in M. L. Lister et al. 2019). It consists of blazars at $\delta>-30\degree$ with VLBI flux density exceeding 1.5 Jy at 15 GHz at least once during the quarter century of observations, 1994–2019. This sample contains 232 blazars, and 205 of those have Doppler boosting measurements (highlighted in Figure 1). The MOJAVE 1.5JyQC sample was specifically designed to be homogeneous and unbiased.
@@ 2.2 IceCube high-energy neutrinos
The first catalog is the sample of high-energy alert-like neutrino detections we collected before, most recently in P23. Earlier, the same selection criteria were utilized in P20 and T. Hovatta et al. (2021). This sample consists of 71 events with the angular uncertainty region below 10 sq. deg. and the estimated energy above 200 TeV.
@@
The sky localization of IceCube events from both samples is shown in Figure 1. Blazars coincident with IceCube events are highlighted in the figure, and those from the MOJAVE 1.5JyQC sample are labeled. See Table 1 for the properties of neutrino-coincident blazars from the full MOJAVE sample. Note how the change of IceCube event parameters between different reconstructions leads to slightly different lists of events and of their blazar coincidences. Evaluating the reconstruction differences in details is out of scope of this paper; here, we use these two catalogs as an extra robustness check of our results, reporting only conclusions consistent between the two.
For the analyses performed in this work, we enlarge the original 90% directional uncertainties reported in IceCube catalogs by $0.78\degree$ isotropically in all directions. This enlargement magnitude was determined in P23 to be optimal for association searches, and we keep the same value to avoid multiple trials in the statistical analysis. As a guidance, Table 1 marks the three coincidences that don’t require this enlargement. A similar enlargement was found in P. M. Kouch et al. (2024); see that paper, P23, and Section 3 for a more detailed discussion.
@@ 3 Neutrino-blazar correlation
2015-09-26b IC150926Ab J1256-0547 1253-055 $0.536$ $15$ 1919 yes $140$ $72$ $0.12$
2022-03-03b J1751++0939 1749++096 $0.322$ $2.8$ 44 yes $100$ $51$ $0.07$
2019-07-30a IC190730Aa J1504++1029 1502++106 $1.838$ $1.5$ $1.4$ yes $65$ $35$ $0.44$
2016-01-28b IC160128Ab J1733-1304 1730-130 $0.902$ 44 $4.5$ yes $48$ $32$ 11
2015-08-12b IC150812Bb J2148++0657 2145++067 $0.999$ $6.6$ $7.5$ yes $31$ $16$ $0.36$
2015-08-31b IC150831Ab J0336++3218 0333++321 $1.259$ $1.6$ $1.7$ yes $30$ $18$ $1.4$
2012-05-15b IC120515Ab J1310++3220 1308++326 $0.997$ $1.9$ $2.1$ yes $24$ $28$ $2.3$
2022-02-05b IC220205Bb J1743-0350 1741-038 $1.054$ 44 $5.3$ yes $22$
2011-09-30b J1743-0350 1741-038 $1.054$ 44 $5.3$ yes $22$
2013-10-14b IC131014Ab J0211++1051 0208++106 $0.2$ $0.66$ $1.1$ yes $6.3$ $5.1$ $8.8$
2020-11-30a IC201130Aa J0206-1150 0203-120 $1.663$ $0.28$ $0.3$ no $35$ 2020 11
2017-09-22b IC170922Ab J0505++0459 0502++049 $0.954$ $0.7$ $0.93$ no $25$ $14$ $1.6$
IC221223Ab J2311++3425 2308++341 $1.817$ $0.69$ $0.86$ no $23$
IC201221Ab J1724++4004 1722++401 $1.049$ $0.5$ $0.58$ no $9.8$ $61$ $3.2$
2010-10-09b J2200++1030 2157++102 $0.172$ $0.19$ $0.22$ no 55
2017-09-22a IC170922Aa J0509++0541 0506++056 $0.3365$ $0.42$ $0.46$ no $1.8$ $1.5$ $33$
2011-07-14b IC110714Ab J0432++4138 0429++415 $1.022$ $1.4$ 11 no $0.13$ $54$ $31$
Note. — Columns are as follows: (1) — IceCube neutrino arrival date for events included in the P23 sample; (2) — IceCat event ID for events included in IceCat-1; see Section 2.2 for a definition of these samples; a indicates neutrino-blazar matches within original reported uncertainties, and b those matches that require extending original uncertainties, following Section 3.
(3), (4) — coordinate-based blazar names, with the J2000 name taken directly from the RFC (L. Y. Petrov & Y. Y. Kovalev, 2025); (5) — redshift $z$ extracted from the NASA/IPAC Extragalactic Database (NED); (6), (7) — parsec-scale radio flux densities measured by VLBI, as presented in the RFC catalog; their typical uncertainty is about $10\%$, dominated by the amplitude calibration; (8) — flag indicating whether the blazar is included in the uniformly-selected MOJAVE 1.5JyQC sample (Section 2.1); those sources are also labeled in Figure 1, and their VLBI images are shown in Figure A1; (9), (10), (11) — beaming indicators, Doppler $D$ and Lorentz $\Gamma$ factors and the viewing angle $\theta$, as measured within the MOJAVE program (D. C. Homan et al., 2021); see Section 4.1 for a discussion of assumptions and uncertainties of these measurements.
@@
After having chosen the test statistic, we calculate the probability that its observed value could have arisen by chance, assuming no directional neutrino-blazar correlation. For that, we randomize the Right Ascensions of all IceCube events $10^{6}$ times, motivated by the instrument sensitivity being dependent on the Declination only (used in M. G. Aartsen et al., 2017, and elsewhere). Then, the chance probability is the fraction of random realizations where the average VLBI flux of matching blazars is higher than the observed value. There are no free parameters to be optimized in this procedure, and we can immediately interpret the resulting fraction as the $p$-value.
@@
Note that updated or more detailed event reconstructions are not necessarily better in every aspect: see demonstrations and discussion in the review S. Troitsky (2024) and in G. Sommani et al. (2024). For example, a reconstruction update lead to a decrease of the TXS 0506+056 flare significance (IceCube Collaboration et al., 2021; R. Abbasi et al., 2023). A detailed analysis and interpretation of the differences between various IceCube event reconstructions is out of scope of this work.
@@
As detailed in Section 2.2, we enlarge original IceCube event uncertainties by $0.78\degree$ for all calculations in this paper, following the approach of P23 and P. M. Kouch et al. (2024) and avoiding multiple comparisons issues. Here, we present a focused cross-check that such an enlargement remains important for both IceCube event catalogs we use (Section 2.2).
Repeating our statistical testing procedure with original IceCube uncertainties taken as-is, we don’t see any significant evidence for the correlation with VLBI blazars. The chance probability $p>5\%$ for both IceCube event selections. The spatial correlation shown above is heavily driven by blazars that lie slightly outside the original reported 90% directional errors. This is still consistent with the observed correlation caused by a physical connection, see P. M. Kouch et al. (2025) for a sensitivity and robustness analysis.
Such an effect could result from unaccounted systematic uncertainties or other discrepancies between estimated and actual error regions. A detailed dedicated study can only be performed by the IceCube collaboration, utilizing all the available data about each detection; see C. Lagunas Gualda et al. (2021) for a discussion and an overview of challenges. Further in the future, it will be crucial to compare the brightest neutrino sources in the sky between different neutrino observatories. The cross-check we performed here does not replace the detailed analysis performed in P23 and P. M. Kouch et al. (2024), but remains consistent with those works and results reported by R. Abbasi et al. (2023, 2024a); C. Bellenghi et al. (2023); J. Kuhlmann & F. Capel (2025).
@@ 4.1 Observational evidence for neutrino beaming
The association of high-energy neutrinos with VLBI-bright blazars (Section 3) already suggests that beaming plays a major role in neutrino production and detection: neutrinos are predominantly emitted along the direction of the jet. Indeed, VLBI-selected AGNs are dominated by galaxies with jets pointing towards us, their typical viewing angles are $\theta=2\dots 3\degree$ (M. L. Lister et al., 2019; D. C. Homan et al., 2021). If neutrino emission were isotropic, we would expect to see thousands ($4\pi/\pi\theta^{2}~\sim~3000$) of similar AGNs pointed in other directions for each neutrino-VLBI association. This is clearly incompatible with the observed number of high-energy neutrinos.
@@
They are measured from VLBI observations within the MOJAVE program (Section 2.1, D. C. Homan et al. 2021). These estimates crucially rely on the VLBI ability to resolve the parsec-scale structure and determine the brightness of the most compact detail – the so-called core. Due to the measurement procedure, the estimates of these quantities are somewhat related to the VLBI flux density we use in the correlation analysis above (Section 3). Still, they are not exactly the same and differ in fundamental ways. The VLBI flux density is a measure of the total radio emission on multi-parsec scale, while Doppler factors are determined from the brightness of even more compact regions, typically a fraction of a parsec in projection. Viewing angle and Lorentz factor estimates additionally rely on parsec-scale kinematics to constrain the geometry. It can be challenging to calculate precise uncertainties of these measurements. Still, as a cross-check, Doppler factors resulting from MOJAVE measurements were found to be consistent with independent estimates obtained via different techniques, such as radio variability (I. Liodakis et al., 2018; D. C. Homan et al., 2021).
@@
We evaluate the statistical significance of the Doppler factor correlation with neutrino coincidences. Doppler factors are chosen because they are the most direct tracers of the electromagnetic emission beaming. They are also measured more directly from observations, see D. C. Homan et al. (2021). We rely on the same statistical testing procedure as in Section 3, namely the shuffling of neutrino event positions in the sky. We choose the Kolmogorov-Smirnov test statistic between the Doppler factor distributions for MOJAVE 1.5JyQC blazars and for neutrino coincidences.
@@
The most extreme viewing angle estimates from the MOJAVE observations can be less than $0.1\degree$. There are only three sources with such viewing angles in the full sample (D. C. Homan et al., 2021), and one of them is coincident with an IceCube neutrino: 1749+096 with $\theta=0.07\degree$, see Table 1 and Figure A1. We believe that such blazars do indeed have jets pointing almost directly towards us, but recommend treating specific values of these extreme viewing angle estimates with caution. It is possible that some limitations of the generic approach utilized in D. C. Homan et al. (2021) manifest in such extreme cases. Additional or alternative ways to infer the jet geometry and the beaming patterns would potentially be useful for more detailed quantitative studies in the future. For example, a very large apparent jet opening angle (A. B. Pushkarev et al., 2017) might be used as another independent indicator of a very small viewing angle (e.g., the case of PKS 1424+240, Y. Y. Kovalev et al. 2025).